### v3

@@ 6. Cost Analysis
Cost of sampling in mini-batch training. Mini-batch training requires two pre-processing steps: sampling and data loading. Data loading costs are included in the communication costs. Modeling sampling costs in terms of FLOPS, as we did for the training, costs $\Theta_{\textrm{mb}}$ is challenging because sampling is a sparse and irregular computation, not a dense computation like matrix multiplication. Therefore, to put the previous results in context, it is necessary to consider the relative computational cost of sampling over training in mini-batch training. We measure it experimentally using DGL and Neighborhood Sampling, running both sampling and training in CPU memory to factor out the overhead of transferring data to the GPUs. The results are shown in Figure 9. The computational cost of sampling is noticeable but is still lower than the training cost. It decreases when we increase the number of workers/partitions because each epoch requires fewer iterations.
Takeaway. This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training, especially with larger graphs. In practice, mini-batch training can replicate input features across multiple workers to eliminate or reduce communication costs. The high communication cost of full-graph training has been the focus of prior work on full-graph training systems, as discussed in Section 2.2.
The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems, especially with larger graphs. Overcoming this gap while maintaining the full-graph training approach and avoiding sampling is challenging.
@@ 7. Conclusion
A common rationale for using full-graph training is to achieve high accuracy by avoiding the information loss results from sampling. In our evaluation, we did not observe these expected benefits. Mini-batch training systems not only achieve a lower time-to-accuracy across all models, datasets, and hardware configurations we consider but also a comparable or even higher accuracy with proper hyperparameter tuning.
Mini-batch training also has some limitations. The common Neighborhood Sampling algorithm is able to consistently achieve good accuracy, but it still shows a slightly larger variance between training runs than vanilla full-graph training. Other sampling algorithms we evaluated are not always able to converge to a high accuracy, depending on the dataset. Finding sampling algorithms that have consistently higher performance than Neighborhood Sampling is an interesting avenue of research. Our experiments show that as GNN models become deeper, the advantage of mini-batch training over full-graph training reduces. Distributed sampling over multiple hosts can also become a bottleneck, especially when the network is slow.
An important advantage of mini-batch training systems is that the sampling algorithms are interchangeable components. This modularity and flexibility in choosing the right sampling algorithm is key since different algorithms perform very differently on different datasets. Our experiments show that providing efficient sampling implementations is important for multi-host setups, especially when using a slow network. In order to support future innovations in sampling algorithms, it is important to have systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021).
In terms of improving full-graph training systems, our cost model analysis and single-GPU performance experiments reveal that in addition to the well-known communication cost of full-graph training—a bottleneck addressed by much previous work—there is a computation cost gap that is difficult to bridge without employing sampling. A natural open question is whether combining full-batch training with more aggressive sampling, forgoing full-graph aggregation, can be more effective than standard mini-batch training. In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantages of using full-batch training are unclear.

### v4

@@ 6. Cost Analysis
Cost of sampling in mini-batch training. Mini-batch training requires two pre-processing steps: sampling and data loading. Data loading costs are included in the communication costs. Modeling sampling costs in terms of FLOPS, as we did for the training, costs $\Theta_{\textrm{mb}}$ is challenging because sampling is a sparse and irregular computation, not a dense computation like matrix multiplication. Therefore, to put the previous results in context, it is necessary to consider the relative computational cost of sampling over training in mini-batch training. We measure it experimentally using DGL and Neighborhood Sampling, running both sampling and training in CPU memory to factor out the overhead of transferring data to the GPUs. The results are shown in Figure 9. The computational cost of sampling is noticeable but is still lower than the training cost in most cases. It decreases when we increase the number of workers/partitions because each epoch requires fewer iterations.
Takeaway. This analysis shows that vanilla full-graph training has a higher communication cost than vanilla mini-batch training when the GNN model is not very deep. In practice, mini-batch training can replicate input features across multiple workers to eliminate or reduce communication costs. The high communication cost of full-graph training has been the focus of prior work on full-graph training systems, as discussed in Section 2.2.
The analysis shows that using sampling in mini-batch training also results in a substantial reduction in computation cost compared to full-graph training systems.
@@ 7. Conclusion
Our work has evaluated various full-graph and mini-batch systems, encompassing a wide range of optimizations (see Figure 2). A common rationale for using full-graph instead of mini-batch training is to achieve higher accuracy by avoiding the information loss resulting from sampling. In our evaluation, we did not observe these expected benefits. Mini-batch training systems not only achieve a lower time-to-accuracy across all models, datasets, and hardware configurations we consider but also a comparable or even higher accuracy with proper hyperparameter tuning.
Among the optimizations we considered for mini-batch training, GPU-based sampling and GPU caching should be used whenever feasible. The choice and configuration of the sampling algorithm is critical for performance and accuracy. The common Neighborhood Sampling algorithm is able to consistently achieve good accuracy, but it still shows a slightly larger variance between training runs than vanilla full-graph training. Other sampling algorithms we evaluated are not always able to converge to a high accuracy, depending on the dataset. Distributed sampling over multiple hosts can also become a bottleneck, especially when the network is slow. Systems that allow ML practitioners to implement new sampling algorithms and run them on GPUs, such as (Jangda et al., 2021; Pandey et al., 2020; Gong et al., 2023; Wang et al., 2021), can support future innovations in efficient sampling algorithms.
For full-graph training systems, the algorithmic optimizations we considered can substantially improve performance with minimal or no impact on accuracy. Our performance evaluation and cost model analysis reveal that in addition to the well-known communication bottleneck addressed by much previous work, future optimizations should also reduce the computation cost gap with mini-batch training. A natural open question is whether combining the full-batch training pipeline of Figure 1(a) with more aggressive sampling, forgoing full-graph aggregation, can be more effective than mini-batch training. In our hyperparameter tuning, however, we found that the accuracy of mini-batching typically peaks with batch sizes around 1024/4096 and does not grow further with larger values, so the advantage of using full-batch training is unclear.